Miss Sun Tzi obtained a level principal and interest on the balance loan of $78,000. The nominal rate of interest is 6% with a term of 5 years and payments scheduled at the end of each month. Compute the 15th payment.

Answers: 1559, 1579, 1599, 1619
Answer #1

Type of loan = level principal payment and interest on balance amount.

Loan amount (PV) = $78000

No of installments (N) = 5 x 12 = 60

Therefore, principal repayment per installment = 78000 / 60 = $1300

Therefore principal repaid in 14 installments = $1300 x 14 = $18,200

Therefore principal outsanding at the end of 14th installment = 78000 - 18200 = $ 59800

Rate of interest (I) = 6% p.a. = 0.50% p.m.

Interest amount for 15th installment = $59800 x 0.50% = $299

Therefore amount of 15th installment = principal payment + interest payment = $1300 + $299 = $1599

Therefore amount of 15th installment = $1599

Thumbs up please if satisfied. Thanks :)